Which term of the geometric sequence is 1/8 if the first term is 4 and common ratio ½


1
Expert's answer
2021-07-06T03:36:06-0400

Solution.

We have geometric sequence, where "b_1=4," "r=\frac{1}{2}" and "b_n=\frac{1}{8}." Find "n."

By the formula "b_n=b_1r^{n-1}" we will have "\frac{1}{8}=4•(\frac{1}{2})^{n-1}."

Solve this equation


"\frac{1}{32}=(\frac{1}{2})^{n-1},\newline
(\frac{1}{2})^5=(\frac{1}{2})^{n-1},\newline
5=n-1,\newline
n=6."

Answer. The sixth term of the geometric sequence is "\frac{1}{8}."